Calculate the voulume occupied by .845 moles of nitrogen gas at pressure of 1.37 atm and a temperature of 315K ? Atm = atmospheres
K= kelvins = temperature
1
Expert's answer
2012-11-12T10:35:45-0500
pV = nRT
p = 1.37atm
n = 0.845 moles
T = 315 K
R = 0.082 L*atm/(K*mol)
V = nRT/p = 0.845*0.082*315/1.37 = 15.93 L